OpenStudy (anonymous):

help! Describe the transformations required to obtain the graph of the function f(x) from the graph of the function g(x). f(x)=9cosx g(x)=cosx a. Vertical stretch by a factor of 9 c. Vertical shrink by a factor of 1/9 b. Horizontal stretch by a factor of 9 d. Horizontal shrink by a factor of 1/9 @Catlover5925 @jordanloveangel

OpenStudy (jordanloveangel):

1/9 f(x) = g(x) thats a stretch factor (1 / 1/9) = 9 parallel to the y-axis

OpenStudy (jordanloveangel):

f(x) is nine times bigger than g(x), so f(x)'s line is 9 times higher than g(x).